OpenStudy (anonymous):

Given that log 6 base 3 is equal to M and log 5 base 6 is equal to N.express log 10 base 3 in terms of M and N

OpenStudy (ash2326):

5 min , i'm solving

OpenStudy (anonymous):

O RIGHT AM WAITING

OpenStudy (ash2326):

M=log 6(3) means log 6 to base 3 N=log 5(6) M=log2(3)+log3(3) M=1+log2(3) so log 2(3)=M-1 N=log5(6) N=log5/log 6 N=(log 5/log 3)/(log 6/log 3) N=log5(3)/M so log 5(3)=MN log10(3)=log 5(3)+log 2(3) log 10(3)=MN+M-1